# Bipartite Fluctuations of Critical Fermi Surfaces

### v2

@@ V Continuous Mott Transition
[…]
@@ V.1 Charge Fluctuations
where each electron $c_{\sigma}$ is fractionalized into a spinless bosonic chargon $b$, which carries the electric charge, and a charge-neutral fermionic spinon $f_{\sigma}$, which carries the spin quantum number. There is a dynamical $\textrm{U}(1)$ gauge field $a_{\mu}=(a_{\tau},\bm{a})$ that couples $b$ and $f$. After introducing the dual bosonic vortices $\varphi$ of the chargons $b$, the critical theory is described by Eq. (19), where $\mathcal{L}_{\textrm{FS}}[f_{\sigma},a]$ represents the Fermi-surface state of spinons. The continuous Mott transition is triggered by the superfluid to Mott insulator transition of chargons. In the dual picture, the vortex condensation is described by
@@
where $\textrm{D}_{\tilde{a}}=\partial-\mathtt{i}\tilde{a}$ denotes the gauge covariant derivative. Here, a single parameter $r\sim(g-g_{c})$ controls the transition, with $g$ denoting the bandwidth of the original electrons. When $r>0$, the vortices $\varphi$ are trivially gapped, resulting in the Landau FL phase of electrons. Conversely, when $r<0$, the condensation of vortices Higgs the gauge field $\tilde{a}$, leading the system to become a spin-liquid Mott insulator with gapless spinons $f_{\sigma}$.
@@
An important feature of the critical theory is that the chargon/vortex sector is dynamically decoupled from the spinon sector, as the conditions outlined at the beginning of Sec. IV.2.2 are indeed satisfied by the XY fixed point [50, 51]. As a result, the zero-temperature response function $\Pi_{b}$ of the chargons follows the CFT form given by Eq. (34), where $\sigma_{xx}^{b}$ represents the critical conductivity of the 3D XY universality class, and the Hall conductivity $\sigma_{xy}^{b}$ vanishes. In the dual picture, the response function of vortices $\Pi_{v}$ is also described by Eq. (34) with the coefficients $\sigma_{xx}^{v}=1/(4\pi^{2}\sigma_{xx}^{b})$ and $\sigma_{xy}^{v}=0$. Assuming the $f$-fermions remain in the mean-field state with the response function $\Pi_{f}$ as given by Eq. (30), we can apply the Ioffe-Larkin rule Eq. (23) (or Eq. (80)) to derive the leading-order long-wavelength behavior of the static structure factor
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ $=\int\frac{\textrm{d}\omega}{2\pi}\frac{1}{\Pi_{f}^{\tau\tau}(k)^{-1}-\frac{\omega}{|\bm{k}|^{2}}\mathsf{F}^{b}_{L}(\frac{\omega}{|\bm{k}|})^{-1}}$
$=-\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots$ (41)
@@
Since the Hall response vanishes, Eq. (41) for the Mott transition is much simpler than Eq. (37) for the CFL-FL transition. The common feature of these two types of critical points is that they correspond to incompressible states with vanishing Drude weight, where the static structure factor exhibits a CFT-like behavior, with the overall coefficient determined by the universal conductivity $\sigma_{xx}^{b}$ of the chargons.
In this case, the bipartite charge fluctuations are described by Eq. (16) for the configuration in Fig. 1, with a universal corner contribution proportional to
$C_{\rho}=C_{J}^{(\textrm{XY})}$ (42)
@@
where $C_{J}^{(\textrm{XY})}=2\sigma_{xx}^{b}/\pi^{2}$ represents the current central charge of the 3D XY fixed point. A notable property arising from time-reversal invariance is the direct relationship between the universal corner contribution and the longitudinal resistivity jump at the critical point. With a vanishing Hall response, the total longitudinal resistivity is given by the Ioffe-Larkin rule $\rho_{xx}=\rho_{xx}^{f}+\rho_{xx}^{b}$, where $\rho_{xx}^{f}$ is the spinon contribution and $\rho_{xx}^{b}$ is the chargon contribution. At $T=0$, in the insulating phase, the gapped chargon sector leads to $\rho_{xx}^{b}=+\infty$, resulting in infinite total electrical resistivity $\rho_{xx}$. In the metallic phase, the chargon superfluid has $\rho_{xx}^{b}=0$, so the total resistivity $\rho_{xx}$ is determined solely by the spinon contribution. At the critical point, a universal jump $\Delta\rho_{xx}(\omega/T)$ is predicted [51] due to the universal chargon contribution. Under the limit $\omega/T\rightarrow\infty$, $\Delta\rho_{xx}=2/(\pi^{2}C_{J}^{(\textrm{XY})})$ is directly related to the universal corner contribution to bipartite charge fluctuations. In the absence of disorder, such that the spinon contribution $\rho_{xx}^{f}\approx 0$, the total critical resistivity can be approximated as $\Delta\rho_{xx}$.
@@ V.2 Spin Fluctuations
Here, the background fields $A^{c}$ and $A^{s}$ are introduced to keep track of the two $\textrm{U}(1)$ symmetries. The spinon $f=(f_{\uparrow},f_{\downarrow})$ still forms a Fermi-surface state and is coupled to emergent $\textrm{U}(1)$ gauge fields $a^{c}$ and $a^{s}$ through the covariant derivative $\textrm{D}=\partial\sigma^{0}-\mathtt{i}a^{c}\sigma^{0}-\mathtt{i}a^{s}\sigma^{3}$, where $\sigma^{\mu}$ denotes the vector of Pauli matrices. The fluxes of the $\textrm{U}(1)$ gauge fields $\tilde{a}^{c}$ and $\tilde{a}^{s}$ represent the densities of the bosonic partons that carry the charge-$\textrm{U}(1)$ and spin-$\textrm{U}(1)$ symmetries. The parton construction for Eq. (43) resembles that of the $\textrm{U}(2)$ gauge theory discussed in Ref. [91]. There is no prior relationship between the two sets of coupling constants $e,r,u$ for the different vortices.
@@
The vortex theory described by Eq. (43) is a multi-critical theory governed by two coupling constants, $r_{c}$ and $r_{s}$. The Mott transition discussed earlier corresponds to the scenario where the spin vortex $\varphi^{s}$ is gapped, while the condensation of the charge vortex $\varphi^{c}$ is dual to the XY transition. When both $\varphi^{c}$ and $\varphi^{s}$ are gapped, the system remains in the ordinary metallic phase. However, if $\varphi^{c}$ condenses, we are left with gapless spinons coupled to a deconfined gauge field $a^{c}$, while the gauge field $\tilde{a}^{c}$ is gapped out due to the Higgs mechanism. This is the Mott insulator with a spinon Fermi surface. Conversely, if $\varphi^{c}$ is gapped and $\varphi^{s}$ condenses, it leads to a transition from an ordinary metal to an algebraic charge liquid with power-law charge correlations. The behaviors of the charge and spin bipartite fluctuations are summarized in Tab. 4. Regarding the final fate in the IR, there is an important difference between the vortex-condensed phases of $\varphi^{c}$ and $\varphi^{s}$. In the case of $\varphi^{s}$, the gauge field $a^{s}$ will eventually induce a pairing instability in the Fermi-surface state, resulting in a ground state of a charge superconductor. Nevertheless, the charge liquid should still be observable within a finite energy window.
@@
Table 4: Leading-order scalings of charge and spin bipartite fluctuations $(\mathcal{F}^{c},\mathcal{F}^{s})$ across the transitions described by the multicritical vortex theory Eq. (43) involving charge and spin vortices $\varphi^{c}$ and $\varphi^{s}$. At the critical point for each vortex condensation, the corresponding bipartite fluctuations exhibit a universal corner contribution proportional to the critical conductivity associated with the 3D XY universality class.
@@ V.3 Charge Fractionalization
The condensation of both $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$ is controlled by a single coupling constant $r$, which is related to the bandwidth of the electrons. In the Mott insulator, each spin/valley sector forms a $\mathbb{Z}_{N}$ topological order. By following a similar analysis to that at the beginning of Sec. IV.2.2, it can be shown that the Landau damping effects are irrelevant at the 3D XY* transition. This implies that, once again, the vortex sector of each spin/valley is dynamically decoupled from the spinon Fermi surface. In both the Mott insulator and the critical point, the charge carriers are the anyons of the $\mathbb{Z}_{N}$ topological order. Each one carries the fractional electric charge $e_{*}=e/N$.
@@
$\mathcal{W}_{\mathcal{C}}^{c}(\vartheta)$ $=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}(\tilde{a}^{\uparrow}+\tilde{a}^{\downarrow})\right),$
$\mathcal{W}_{\mathcal{C}}^{s}(\vartheta)$ $=\exp\left(\frac{\mathtt{i}\vartheta}{2\pi}\int_{\mathcal{C}}(\tilde{a}^{\uparrow}-\tilde{a}^{\downarrow})\right),$ (48)
which define the charge and spin bipartite fluctuations $\mathcal{F}_{\mathcal{A}}^{c}$ and $\mathcal{F}_{\mathcal{A}}^{s}$ through Eq. (25), where $\mathcal{C}=\partial\mathcal{A}$. The scaling behaviors of $\mathcal{F}^{c}$ can be easily seen to follow $L\log L$ and $L$ by using the gauge-field propagators $\langle\tilde{a}^{\uparrow}\tilde{a}^{\uparrow}\rangle$ and $\langle\tilde{a}^{\downarrow}\tilde{a}^{\downarrow}\rangle$ before and after the vortex condensation. At the critical point, an interesting universal corner contribution appears, with a coefficient that deviates from Eq. (42). The self-energy of $\tilde{a}^{\uparrow}$ is dominated by the loop corrections from $\varphi^{\uparrow}$, which is proportional to $N^{2}$. The two-point function of the conserved current $J^{\uparrow}=\frac{\mathtt{i}}{2\pi}\star\textrm{d}\tilde{a}^{\uparrow}$ at the XY* fixed point can be shown to be proportional to $1/N^{2}$ and follows the form
@@
where we still use $C_{J}^{(\textrm{XY})}$ to denote the current central charge at the XY fixed point. Identical expressions hold for $\tilde{a}^{\downarrow}$ and $J^{\downarrow}$. As a result, the charge fluctuations $\mathcal{F}^{c}$ now follow Eq. (16) with a universal coefficient
@@
This value of $C_{\rho}$ can be used to determine the universal resistivity jump $\Delta\rho=N^{2}/(\pi^{2}C_{J}^{(\textrm{XY})})$, as delineated in Eq. (15). In Ref. [52], it is argued that disorder effects are weak. Neglecting the contributions to the total resistivity from the spinon Fermi surface, we approximate the total resistivity by the jump $\Delta\rho$. The factor of $N^{2}$ enhancement in $\Delta\rho$ may potentially explain the large critical resistivity observed in Ref. [52].
Another key difference compared to the Mott transition described by Eq. (43) is the distinct scaling of the bipartite spin fluctuations $\mathcal{F}^{s}$ in the Mott insulator. Because the vortex condensation Higgses both $\tilde{a}^{\uparrow}$ and $\tilde{a}^{\downarrow}$, the scaling follows a boundary law $\mathcal{F}^{s}\sim L$.
@@ VI Discussion and Outlook
In this paper, we investigate the bipartite fluctuations of conserved quantities across interaction-driven “metal-insulator transitions” from a Landau FL to a “charge insulator” defined by localization criteria (see Sec. II.2). These charge insulators, which exhibit a vanishing Drude weight, include incompressible examples like spin-liquid Mott insulators and compressible examples like CFLs. If the transitions are continuous, we predict a logarithmic divergence of the localization tensor as the correlation length approaches infinity, as described in Eq. (14). This divergence is associated with a universal subleading corner contribution to bipartite fluctuations (see Eq. (16)), with the universal coefficient determined by the critical resistivity jump, as outlined in Eq. (15). Therefore, we establish a fluctuation-dissipation theorem-type relationship linking ground-state charge (or spin) fluctuations to transport properties. Our results, which stem from the general consequences of the Ioffe-Larkin composition rule, are based on two dynamically decoupled parton sectors at criticality: one forms a stable Fermi-surface state, while the other behaves like a CFT at the transition. These results are expected to hold at other similar phase transitions as well.
In Sec. IV, we also examine the CFL phase, where the localization tensor $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ is finite, and a subleading corner contribution is identified in Eq. (12). Based on the standard RPA analysis of the HLR theory, we find that the value of $\mathsf{G}$ is universal and is given by the Hall conductivity $\sigma_{xy}$ (see Eq. (13)). However, as we will clarify in an upcoming work [75], for a general (anisotropic) charge insulator, the value of $\mathsf{G}^{ij}$ is non-universal and is subject to both a topological lower bound and an energetic upper bound [60, 61, 62, 72]
00 0 $\leq\mathsf{G}^{ij}-\frac{\mathtt{i}}{2}\sigma_{xy}\epsilon^{ij}$
00 0 $\leq\lim_{\bm{A}\rightarrow 0}\frac{1}{2\Delta}\frac{1}{V}\left\langle\!\frac{\partial^{2}H}{\partial A_{i}\partial A_{j}}\!\right\rangle-\mathsf{G}^{ij}$ (51)
where the right-hand side of each line is a positive semi-definite matrix. Here, $\Delta$ denotes the energy gap of the many-body Hamiltonian $H$, $V$ is the volume of the system, and $\bm{A}$ represents the background U(1) field. The specific conditions for the saturation of these bounds will be further elaborated in Ref. [75].

### v3

@@ V Continuous Mott Transition
In this section, we now turn to spin-$1/2$ electrons, which enable the definition of both charge and spin fluctuations within a subsystem. We examine two theoretical constructions of continuous Mott transitions at half-filling that preserve time-reversal symmetry. Due to their technical similarities to the CFL-FL transition, some of the results from Sec. IV are applicable here. We begin with the original proposal [50, 51, 52], which was motivated by the Mott organic compound $\kappa$-$\textrm{(ET)}_{2}\textrm{Cu}_{2}\textrm{(CN)}_{3}$ [102], and also discuss a generalized multicritical vortex theory in this context. We then turn to a modified theory [55] proposed for another candidate material, the TMD morié bilayer $\textrm{MoTe}_{2}/\textrm{WSe}_{2}$, which exhibits an anomalously large critical resistivity in the experiment [53].
@@ V.1 Charge Fluctuations
where each electron $c_{\sigma}$ is fractionalized into a spinless bosonic chargon $b$, which carries the electric charge, and a charge-neutral fermionic spinon $f_{\sigma}$, which carries the spin quantum number $\sigma=\uparrow,\downarrow$. There is a dynamical $\textrm{U}(1)$ gauge field $a_{\mu}$ that couples $b$ and $f$. After introducing the dual bosonic vortices $\varphi$ of the chargons $b$, the critical theory is described by Eq. (19), where $\mathcal{L}_{\textrm{FS}}[f_{\sigma},a]$ describes a Fermi surface of spinons. The continuous Mott transition is driven by a superfluid-to-Mott-insulator transition of the chargons. In the dual description, this is captured by the condensation of vortices, governed by
@@
where $\textrm{D}_{\tilde{a}}=\partial-\mathtt{i}\tilde{a}$ is the gauge covariant derivative. A single tuning parameter $r\sim(g-g_{c})$, with $g$ denoting the electronic bandwidth, controls the transition. For $r>0$, the vortex field $\varphi$ is trivially gapped, corresponding to the Landau FL phase of electrons. For $r<0$, vortex condensation Higgses the gauge field $\tilde{a}$, driving the system into a spin-liquid Mott insulator characterized by a gapless spinon Fermi surface. The expectation for the distinct scalings of bipartite fluctuations in the two phases was briefly mentioned in Ref. [103].
@@
An important feature of the critical theory is the dynamical decoupling between the chargon (or vortex) sector and the spinon sector, as the conditions outlined in Sec. III.2 are satisfied by the 3D XY fixed point [51, 52]. As a result, the zero-temperature response function $\Pi_{b}$ of the chargons follows the CFT form in Eq. (37), with $\sigma_{xx}^{b}$ being the critical conductivity of the 3D XY universality class, and $\sigma_{xy}^{b}=0$. In the dual picture, the response function of vortices $\Pi_{v}$ is also of the same form Eq. (37), with the coefficients $\sigma_{xx}^{v}=1/(4\pi^{2}\sigma_{xx}^{b})$ and $\sigma_{xy}^{v}=0$. Assuming the $f$-fermions remain in their mean-field state with response function $\Pi_{f}$ given by Eq. (33), the Ioffe-Larkin rule Eq. (23) or Eq. (88) yields the leading long-wavelength behavior of the static structure factor
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})$ $=\int\frac{\textrm{d}\omega}{2\pi}\frac{1}{\Pi_{f}^{\tau\tau}(k)^{-1}+\frac{\omega}{|\bm{k}|^{2}}\mathsf{F}_{L}^{b}(\frac{\omega}{|\bm{k}|})^{-1}}$
$=\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots$ (44)
@@
Since the Hall response vanishes, the expression for the Mott quantum criticality in Eq. (44) takes a simpler form compared to that for the CFL-FL transition in Eq. (40). A common feature of both transitions is that they describe incompressible states with vanishing Drude weight, characterized by a static structure factor exhibiting CFT-like behavior with an overall coefficient set by the universal conductivity $\sigma_{xx}^{b}$ of the chargons.
For the subsystem geometry shown in Fig. 1, the bipartite charge fluctuations follow Eq. (16), and feature a universal corner contribution proportional to the current central charge of the 3D XY fixed point
$C_{\rho}=\frac{2}{\pi^{2}}\sigma_{xx}^{b}=C_{J}^{(\textrm{XY})}$ (45)
@@
Due to time-reversal invariance, the universal coefficient is directly tied to the longitudinal resistivity jump at the critical point. With a vanishing Hall response, the total longitudinal resistivity is given by the Ioffe-Larkin rule $\rho_{xx}=\rho_{xx}^{f}+\rho_{xx}^{b}$, with $\rho_{xx}^{f}$ and $\rho_{xx}^{b}$ denoting the spinon and chargon contributions, respectively. At zero temperature, the insulating phase features $\rho_{xx}^{b}=+\infty$ due to the chargon gap, leading to diverging total resistivity. In contrast, in the metallic phase, the chargon superfluid leads to $\rho_{xx}^{b}=0$, so $\rho_{xx}$ is entirely determined by the spinon sector. At the critical point, the universal chargon contribution induces a finite jump $\Delta\rho_{xx}(\omega/T)$ in the longitudinal resistivity [52]. In the limit $\omega/T\to\infty$, this jump is given by $\Delta\rho_{xx}=2/(\pi^{2}C_{J}^{(\textrm{XY})})$, establishing a direct link between the critical resistivity and the universal corner term in bipartite charge fluctuations. In clean systems with negligible spinon resistivity $\rho_{xx}^{f}\approx 0$, the total critical resistivity is well approximated by $\Delta\rho_{xx}$.
@@ V.2 Spin Fluctuations
Here, the background fields $A^{c}$ and $A^{s}$ are introduced to keep track of the global charge and spin $\textrm{U}(1)$ symmetries, respectively. The spinon field $f=(f_{\uparrow},f_{\downarrow})$ forms a Fermi surface and is minimally coupled to emergent $\textrm{U}(1)$ gauge fields $a^{c}$ and $a^{s}$ through the covariant derivative $\textrm{D}=\partial\sigma^{0}-\mathtt{i}a^{c}\sigma^{0}-\mathtt{i}a^{s}\sigma^{3}$, where $\sigma^{\mu}$ denotes the vector of Pauli matrices. The fluxes of the $\textrm{U}(1)$ gauge fields $\tilde{a}^{c}$ and $\tilde{a}^{s}$ represent the densities of bosonic partons that carry the global charge and spin $\textrm{U}(1)$ quantum numbers, respectively. The parton construction underlying Eq. (46) is reminiscent of the $\textrm{U}(2)$ gauge theory framework discussed in Ref. [104]. There is no a priori relation between the two sets of couplings $e,r,u$ associated with the charge and spin vortex sectors.
@@
The vortex theory described by Eq. (46) is a multi-critical theory governed by two coupling constants, $r_{c}$ and $r_{s}$. The Mott transition discussed earlier corresponds to the situation where the spin vortex $\varphi^{s}$ remains gapped, while the condensation of the charge vortex $\varphi^{c}$ is dual to the 3D XY transition. When both $\varphi^{c}$ and $\varphi^{s}$ are gapped, the system stays in the ordinary metallic phase. In contrast, if $\varphi^{c}$ condenses, the system enters a Mott insulating phase characterized by gapless spinons coupled to a deconfined gauge field $a^{c}$, with $\tilde{a}^{c}$ gapped out by the Higgs mechanism. Conversely, if $\varphi^{c}$ remains gapped while $\varphi^{s}$ condenses, the transition leads to an algebraic charge liquid, featuring power-law charge correlations. The expected behaviors of charge and spin bipartite fluctuations are summarized in Tab. 4. Regarding the final fate in the IR, there is an important distinction between the two phases where a single vortex condenses. In the $\varphi^{s}$-condensed case, the gauge field $a^{s}$ eventually drives a pairing instability of the Fermi-surface state, resulting in a superconducting ground state. Nevertheless, the intermediate charge liquid phase should still be observable within a finite energy window.
@@
Table 4: Leading-order scalings of charge and spin bipartite fluctuations $(\mathcal{F}^{c},\mathcal{F}^{s})$ across the transitions governed by the multicritical vortex theory Eq. (46) involving charge and spin vortices $\varphi^{c}$ and $\varphi^{s}$. At the critical point associated with each vortex condensation, the bipartite fluctuations exhibit a universal corner contribution proportional to the current central charge of the 3D XY fixed point.
@@ V.3 Charge Fractionalization
The condensation of both $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$ is controlled by a single coupling constant $r$, which is related to the bandwidth of the electrons. In the Mott insulator, each spin/valley sector forms a $\mathbb{Z}_{N}$ topological order. Following an analysis similar to that in Sec. III.2, it can be shown that Landau damping effects are irrelevant at the 3D XY* transition. As before, the vortex sector associated with each spin (or valley) is dynamically decoupled from the spinon Fermi surface. At both the Mott insulating phase and the critical point, the charge carriers are anyons associated with the $\mathbb{Z}_{N}$ topological order, each carrying a fractional electric charge of $e_{*}=e/N$.
@@
[…]
which define the charge and spin bipartite fluctuations $\mathcal{F}_{\mathcal{A}}^{c}$ and $\mathcal{F}_{\mathcal{A}}^{s}$ through Eq. (25), where $\mathcal{C}=\partial\mathcal{A}$. The scaling behaviors of $\mathcal{F}^{c}$ can be easily seen to follow $L\log L$ and $L$ by using the gauge-field propagators $\langle\tilde{a}^{\uparrow}\tilde{a}^{\uparrow}\rangle$ and $\langle\tilde{a}^{\downarrow}\tilde{a}^{\downarrow}\rangle$ before and after the vortex condensation. At the critical point, an interesting universal corner contribution appears, with a coefficient that deviates from Eq. (45). The self-energy of $\tilde{a}^{\uparrow}$ is dominated by the loop corrections from $\varphi^{\uparrow}$, which is proportional to $N^{2}$. Consequently, the two-point function of the conserved current $J^{\uparrow}=\frac{\mathtt{i}}{2\pi}\star\textrm{d}\tilde{a}^{\uparrow}$ at the XY* fixed point takes the form
@@
where we still use $C_{J}^{(\textrm{XY})}$ to denote the current central charge at the XY fixed point. Identical expressions hold for $\tilde{a}^{\downarrow}$ and $J^{\downarrow}$. As a result, the charge fluctuations $\mathcal{F}^{c}$ again follow Eq. (16), now with the universal coefficient
@@
This value of $C_{\rho}$ allows one to determine the universal resistivity jump $\Delta\rho_{xx}=N^{2}/(\pi^{2}C_{J}^{(\textrm{XY})})$, as given in Eq. (15). It is argued in Ref. [53] that disorder effects in this system are weak. Neglecting the spinon Fermi surface contribution to the total resistivity, we approximate the full resistivity at criticality by $\Delta\rho_{xx}$. The factor of $N^{2}/2$ enhancement in $\Delta\rho_{xx}$ may offer a possible explanation for the large critical resistivity observed in Ref. [53].
Another key distinction from the Mott transition described by Eq. (46) lies in the scaling behavior of the bipartite spin fluctuations $\mathcal{F}^{s}$ in the Mott insulating phase. Due to the vortex condensation, both $\tilde{a}^{\uparrow}$ and $\tilde{a}^{\downarrow}$ are Higgsed, leading to a boundary-law scaling $\mathcal{F}^{s}\sim L$.
@@ VI Discussion and Outlook
In this paper, we investigate the bipartite fluctuations of conserved quantities across interaction-driven “metal-insulator transitions” from a Landau FL to a “charge insulator” defined by localization criteria (see Sec. II.2). These charge insulators, which exhibit a vanishing Drude weight, include incompressible examples like spin-liquid MIs and compressible ones like CFLs. If the transitions are continuous, we predict a logarithmic divergence of the localization tensor as the correlation length approaches infinity, as described in Eq. (14). This divergence is associated with a universal subleading corner contribution to bipartite fluctuations (see Eq. (16)), with the universal coefficient determined by the critical resistivity jump, as outlined in Eq. (15). Our results reveal a generalized fluctuation-response relation between ground-state charge (or spin) fluctuations and transport properties.
These findings stem from the general structure imposed by the Ioffe-Larkin composition rule and the presence of two dynamically decoupled parton sectors at criticality: one forms a stable Fermi-surface state, while the other behaves like a CFT. These results are expected to hold at other similar phase transitions as well.
Below, we provide further discussions on our central results and related topics while proposing interesting directions for future research.
VI.1 Non-Fermi Liquids